Last year a portfolio manager at PIB bought two million shares of Volkswagen’s stock at €75 per share and sold it one year later at €80. The spot exchange rate one year ago was $1.50 = €1 and the spot rate prevailing at the end of the year was $1.40 = €1.

a. What are the values of portfolio manager’s gain(loss) in Euro (provide amount and percentage)?

b. What the values of portfolio manager’s gain(loss) in USD (provide amount and percentage)?

c. What could possibly explain the difference between manager’s portfolio performance in Euro and USD.
Answer #1

a]

Gain in Euro (Amount) = (selling price - buying price) * number of shares

Gain in Euro (Amount) = (€80 - €75) * 2 million = €10,000,000

Gain in Euro (%) =   (selling price - buying price) / buying price =  (€80 - €75) / €75 = 6.67%

b]

Buying price in USD = Euro buying price * exchange rate 1 year ago = €75 * 1.50 =   $112.50

Selling price in USD = Euro selling price * exchange rate at end of year = €80 * 1.40 =   $112

Gain in USD (Amount) = (selling price - buying price) * number of shares

Gain in USD (Amount) = ($112 - $112.50) * 2 million = -$1,000,000 (loss)

Gain in USD (%) =   (selling price - buying price) / buying price =  ($112 - $112.50) / $112.50 = -0.44% (loss)

c]

The difference is due to the depreciation of the Euro against USD over the 1-year period.